# Enhancing Agentic textual Graph Retrieval with Synthetic Stepwise Supervision

### v1

@@ Abstract
A significant portion of real-world data is inherently represented as textual graphs, and integrating these graphs into large language models (LLMs) is promising to enable complex graph-based question answering. However, a key challenge in LLM-based textual graph QA systems lies in graph retrieval, i.e., how to retrieve relevant content from large graphs that is sufficiently informative while remaining compact for the LLM context. Existing retrievers suffer from poor performance since they either rely on shallow embedding similarity or employ interactive retrieving policies that demand excessive data labeling and training cost. To address these issues, we present Graph-S3, an agentic textual graph reasoning framework that employs an LLM-based retriever trained with synthetic stepwise supervision. Instead of rewarding the agent based on the final answers, which may lead to sparse and unstable training signals, we propose to closely evaluate each step of the retriever based on offline-extracted golden subgraphs. Our main techniques include a data synthesis pipeline to extract the golden subgraphs for reward generation and a two-stage training scheme to learn the interactive graph exploration policy based on the synthesized rewards. Based on extensive experiments on three common datasets in comparison with seven strong baselines, our approach achieves an average improvement of 8.1% in accuracy and 9.7% in F1 score. The advantage is even higher in more complicated multi-hop reasoning tasks. Our code will be open-sourced.
@@ 1 Introduction
Textual graphs are graphs with text-attributed nodes and edges, which are widely used for structured knowledge representation with many applications in question answering, recommendation, and scientific discovery (Peng et al., 2024; Procko & Ochoa, 2024; Zhang et al., 2025a). By explicitly modeling multi-hop relations and semantic constraints, textual graphs enable interpretable and compositional reasoning that is difficult to achieve with unstructured text corpora (Chen et al., 2020; Hogan et al., 2021; Zou, 2020).
The early approaches to reasoning over textual graphs relied on costly annotations and inflexible symbolic inference (Yih et al., 2016). The emergence of large language models (LLMs) has alleviated these limitations through strong semantic understanding (Chang et al., 2024), inspiring a growing body of work that combines LLMs with textual graphs for general-purpose graph understanding and question answering (QA) (Lewis et al., 2020; Peng et al., 2024; Procko & Ochoa, 2024; Zhang et al., 2025a; Jin et al., 2024; Chai et al., 2023).
A key step in LLM-based textual graph QA systems is graph retrieval, i.e., retrieving the relevant content from the target graph based on the natural language query. Most existing retrievers encode graph nodes and edges as vector embeddings and perform similarity-based matching with the query (Chen et al., 2024b; a; Tang et al., 2024). This approach has the advantage of efficiency, but it often produces noisy or incomplete results due to the coarse-grained matching process. Moreover, these methods usually retrieve a large neighborhood in a single step and then flatten all candidate triples into text for the LLM (Edge et al., 2024; Guo et al., 2024b; He et al., 2024). Such flattening discards the relational structure of the graph and obscures reasoning trajectories, which degrades performance on long multi-hop reasoning tasks.
Another line of work employs LLM-based agents that access graph information through tool calls (Jiang et al., 2024; Yang et al., 2024; Ji et al., 2024). Although these methods have the potential to surpass simple similarity retrieval, their training is typically bootstrapped using supervised fine-tuning (SFT). This strategy causes the action space of the model to collapse (Chu et al., 2025; Li et al., 2024), leading it to memorize patterns from the training data rather than learning generalizable policies. As a result, such agents struggle to locate optimal reasoning trajectories. Moreover, constructing these trajectories requires substantial expert annotations, making such approaches difficult to scale.
@@
To this end, we propose Graph-S3, an agentic retrieval framework that equips an LLM-based retriever with the ability to perform interactive, structure-aware exploration of textual graphs. Figure 1 provides an example of agentic retrieval, showing how the retriever agent iteratively performs actions to collect the necessary information to solve the given query.
To train such an LLM-based agent system, a straightforward approach is to employ outcome-supervised reinforcement learning (Lightman et al., 2023; Paolo et al., 2024), i.e., rolling out lots of reasoning trajectories in the graphs with a policy and optimize the policy with reward signals computed from the final answers. However, this approach is difficult to apply in practice for textual graph reasoning, since the outcome-based rewards are often sparse and unstable: The action space for real-world textual graphs is often too large to efficiently discover the optimal retrieval path, while redundant or erroneous retrieval steps may still lead to the correct final answer, making outcome supervision an unreliable signal of reasoning quality (Liu et al., 2023; Rengarajan et al., 2022). This challenge hinders the direct adoption of standard RL algorithms to the problem of textual graph reasoning.
@@
Extensive experiments have demonstrated the effectiveness of our synthetic stepwise supervision approach. For example, while retrieving only 11.44% of the triples, Graph-S3 achieves an average improvement of 8.1% in accuracy and 9.7% in F1 score across the WebQSP, CWQ, and MetaQA benchmarks.
@@
(2) We design a two-stage training paradigm tailored for graph reasoning: supervised fine-tuning on raw synthetic trajectories to bootstrap basic navigation ability, followed by reinforcement learning with synthetic stepwise rewards on refined trajectories to provide explicit feedback and strengthen reasoning strategies.
@@ 2.1 Graph Retrieval Methods
Graph retrieval approaches include similarity-based, GNN-based, and LLM-based methods (Peng et al., 2024; Procko & Ochoa, 2024; Zhang et al., 2025a; Zhu et al., 2025; Han et al., 2024), but most perform one-shot retrieval and often return redundant or incomplete subgraphs. Recent interactive frameworks (Jiang et al., 2024; Ji et al., 2024; Yang et al., 2024) allow iterative exploration, yet their training predominantly relies on imitation of language patterns or outcome-based supervision, which provides only coarse feedback and limits stable multi-hop reasoning. In contrast, our work employs reinforcement learning with synthetic stepwise rewards and a scalable data synthesis pipeline to provide supervision for interactive retrieval.
2.2 Stepwise Reinforcement Learning for Graph Reasoning
Recent advances such as OpenAI o1 and DeepSeek-R1 (Jaech et al., 2024; Guo et al., 2025) demonstrate the effectiveness of reinforcement learning in strengthening multi-step reasoning, enabling models to perform longer chains of thought with improved reliability in domains like mathematics and programming (Guo et al., 2024a; El-Kishky et al., 2025). In contrast, applying RL to textual graphs remains limited, partly due to the lack of fine-grained supervision data (Zhang et al., 2025a; Yao et al., 2025; Liu et al., 2025). RL-based graph agents  (Das et al., 2017; Cui et al., 2025) relied on sparse outcome rewards, making credit assignment across reasoning trajectories difficult. Although recent efforts have introduced reasoning-structured datasets (Pahilajani et al., 2024), high-quality stepwise supervision for graph-based RL remains scarce and difficult to construct at scale. These observations highlight the need for scalable approaches that can provide fine-grained supervision signals and support stable optimization for interactive graph retrieval.
@@ 3 Method
We present our agentic retrieval framework, designed to equip large language models with robust graph reasoning capabilities through stepwise supervision, a two-stage training paradigm, and an interactive retrieval strategy. As illustrated in Figure 2, the framework comprises three main components. First, we construct an automatic data synthesis pipeline that leverages GPT-4o (OpenAI, 2024) to generate diverse exploratory trajectories, which are subsequently refined into high-quality stepwise supervision data. This addresses the scarcity of fine-grained training signals for graph-based reinforcement learning. Second, we adopt a two-stage training paradigm: SFT on raw synthetic trajectories provides a warm-up initialization for basic graph navigation, while RL with synthetic stepwise rewards on refined trajectories supplies explicit feedback at each decision step, stabilizing optimization and strengthening reasoning strategies. Finally, during inference, the retriever operates under an interactive retrieval mechanism that conducts stepwise, structure-aware exploration of the textual graph, thereby reducing redundancy and mitigating incomplete retrieval.
@@ 3.1.1 Action Space for Graph Exploration
$\text{Explore}(x)$ $=\{(x,r,y)\mid(x,r,y)\in\mathcal{G}\}\cup\{(y,r,x)\mid(y,r,x)\in\mathcal{G}\},$ (1)
@@
$\operatorname{Choose}(q,\mathcal{G}^{p})=\{(x,r,y)\in\mathcal{G}^{p}\mid F(q,(x,r,y))=1\}.$ (2)
@@ 3.1.2 Graph Reasoning Data Synthesis
Transition. Executing $a_{t}$ updates the state via graph action space:
$s_{t+1}=\begin{cases}\big(q,\;\mathcal{G}^{p}_{t}\cup\textsc{Explore}(x_{t}),\;\mathcal{G}^{sub}_{t},\;h_{t}\!\cup\!\{a_{t}\}\big),&a_{t}=\textsc{Explore}(x_{t}),\\[2.0pt] \big(q,\;\mathcal{G}^{p}_{t},\;\{(x,r,y)\!\in\!\mathcal{G}^{p}_{t}\;\mid\;F_{B}(q,(x,r,y))=1\},\;h_{t}\!\cup\!\{a_{t}\}\big),&a_{t}=\textsc{Choose},\\[2.0pt] \big(q,\;\mathcal{G}^{p}_{t},\;\mathcal{G}^{sub}_{t},\;h_{t}\!\cup\!\{a_{t}\}\big),&a_{t}=\textsc{Finish},\end{cases}$ (4)
An episode terminates when $a_{t}=\textsc{Finish}$ or when $t$ reaches a preset limit $T_{\max}$.
Answer and retention.
Upon termination at step $T$, we produce an answer $\hat{y}\;=\;\text{Answer}\!\big(q,\mathcal{G}^{sub}_{T}\big)$ and keep the trajectory $\tau\;=\;\big\{(s_{t},a_{t})\big\}_{t=1}^{T}$ only if $\hat{y}$ matches the set of ground-truth answers. After that, we use the raw action labels for SFT dataset: $\mathcal{D}_{\text{SFT}}\;=\;\big\{(s_{t},a_{t})\big\}_{\tau,\,t}$. In the next subsection, we further refine trajectories to obtain stepwise supervision signals for RL.
3.1.3 Trajectory Refinement for Reinforcement Learning
While supervised fine-tuning directly benefits from raw trajectories, reinforcement learning requires more concise training signals (Yue et al., 2025). Filtering trajectories only by final answer correctness often produces redundant exploration steps, introducing noise and inefficiency during policy optimization. To address this, we introduce a refinement procedure that removes unnecessary detours while preserving all answer-consistent trajectories, thereby yielding shorter and cleaner trajectories for RL.
@@ 3.1.3 Trajectory Refinement for Reinforcement Learning
where $\mathcal{F}_{i}$ is the set of all feasible answer-consistent trajectories equivalent to $\tau_{i}$:
$\mathcal{F}_{i}=\left\{\tau\;\middle|\;\text{FinalState}(\tau)=\text{FinalState}(\tau_{i}),\;\text{CorrectAnswer}(\tau)=\text{CorrectAnswer}(\tau_{i})\right\}.$ (7)
Thus, the refined dataset $\mathcal{T}^{*}$ retains trajectories that are semantically equivalent to the originals but stripped of redundant exploration steps. Each refined trajectory $\tau_{i}^{*}$ is then decomposed into step-level supervision signals by attaching a rule-based stepwise reward $\ell_{t}\in[0,1]$ to each action $a_{t}$, indicating its correctness within the reasoning trajectory. Formally, the RL training dataset is constructed as $\mathcal{D}_{\text{RL}}=\big\{(s_{t},a_{t},\ell_{t})\big\}_{\tau^{*},\,t}$. This ensures that reinforcement learning receives concise stepwise supervision signals, improving both stability and efficiency of training.
@@ 3.2.1 Stage I: Supervised fine-tuning
$\mathcal{L}_{\mathrm{SFT}}(\theta)=-\,\mathbb{E}_{(s_{t},y_{t})\sim\mathcal{D}_{\text{SFT}}}\Bigg[\sum_{l=1}^{L_{t}}\log\pi_{\theta}\!\big(y_{t}^{\,l}\,\big|\,\mathcal{I}(s_{t}),\,y_{t}^{<l}\big)\Bigg],$ (8)
@@
3.2.2 Stage II: Reinforcement learning with stepwise rewards
@@
$\ell_{t}=\begin{cases}0,&\text{if }\mathrm{Invalid}(a_{t}),\\[6.0pt] c_{1},&\text{if }\mathrm{Format}(a_{t})=1\;\wedge\;\mathrm{ActionCorrect}(a_{t})=0,\\[6.0pt] c_{2},&\text{if }\mathrm{Partial}(a_{t})=1,\\[6.0pt] 1.0,&\text{if }a_{t}=a_{t}^{*}.\end{cases}$ (9)
Here $\mathrm{Invalid}(\cdot)$, $\mathrm{Format}(\cdot)$, and $\mathrm{Partial}(\cdot)$ are deterministic rule-based functions that check output validity, format correctness, and partial correctness, respectively, and $c_{1},c_{2}$ are dataset-specific hyperparameters. Consequently, we adopt the reward function shown in Eq. 9 to train our model using the GRPO method.
@@ 3.3 Interactive Retriever
At inference time, Graph-S3 interacts with the textual graph through the defined action space. Unlike single-pass retrieval methods that return large subgraphs, our approach performs stepwise exploration by balancing EXPLORE and CHOOSE actions, terminating with FINISH when sufficient evidence is gathered. This interactive process enables precise control over retrieval depth while minimizing redundancy, producing concise subgraphs for reasoning.
@@ 4 Experiments
KG-Agent + Qwen3-8B 29.66 38.02 15.64 22.45 75.80 85.92 28.41 33.66 4.07 10.89
Graph-S3 + Qwen3-8B 36.24 47.88 17.87 23.29 81.50 90.22 53.73 65.60 12.73 29.49
@@
KG-Agent + Llama3.1-8B 32.04 41.99 10.82 13.75 65.55 75.03 32.51 44.52 3.85 8.29
Graph-S3 + Llama3.1-8B 32.31 43.26 17.11 21.17 67.50 76.56 40.17 55.49 10.73 29.55
@@
KG-Agent + Finetuned-8B 42.60 55.38 13.77 16.29 75.99 86.02 30.19 43.95 2.64 7.20
Graph-S3 + Finetuned-8B 44.29 58.45 23.62 30.44 82.77 92.04 63.17 76.18 14.70 36.29

### v3

@@ Abstract
Integrating textual graphs into Large Language Models (LLMs) is promising for complex graph-based QA. However, a key bottleneck is retrieving informative yet compact subgraphs that fit the LLM context. Existing retrievers often struggle, relying either on shallow embedding similarity or costly interactive policies that require excessive supervision. To address these challenges, we introduce Graph-S3, an agentic textual graph reasoning framework featuring an LLM-based retriever trained with synthetic stepwise supervision. Rather than relying on final answer rewards—which often yield sparse and unstable signals—we optimize the retriever by evaluating each step against offline-extracted golden subgraphs. Our approach distills golden subgraphs via a specialized data synthesis pipeline to formulate dense rewards, facilitating a two-stage training scheme that effectively learns the interactive graph exploration policy. Based on extensive experiments on three common datasets in comparison with several strong baselines, our approach achieves an average improvement of 15.6% in accuracy and 17.2% in F1 score. The advantage is even higher in more complicated multi-hop reasoning tasks.
Enhancing Agentic Textual Graph Retrieval with Synthetic Stepwise Supervision
[…]
11footnotetext: Equal contribution.22footnotetext: Corresponding author.33footnotetext: Work done during internship at AIR, Tsinghua University.
@@ 1 Introduction
Textual graphs are widely employed for structured knowledge representation in domains like question answering and scientific discovery (Peng et al., 2024; Procko and Ochoa, 2024; Zhang et al., 2025a). By explicitly modeling multi-hop relations, they enable interpretable, compositional reasoning that is difficult to achieve with unstructured text (Chen et al., 2020; Hogan et al., 2021; Zou, 2020). While early methods were constrained by inflexible symbolic inference and high annotation costs (Yih et al., 2016), the semantic capabilities of Large Language Models (LLMs) have alleviated these limitations (Chang et al., 2024; Su et al., 2025). Consequently, integrating LLMs with textual graphs has become a prominent paradigm for general-purpose graph understanding and QA (Lewis et al., 2020; Peng et al., 2024; Procko and Ochoa, 2024; Zhang et al., 2025a; Jin et al., 2024; Chai et al., 2023).
@@
LLM agents accessing graphs via tool calls surpass simple retrieval (Jiang et al., 2024; Yang et al., 2024; Ji et al., 2024; Xu et al., 2026), as illustrated in Figure 1 where an agent iteratively collects information. However, training such agents via supervised fine-tuning (SFT) often collapses the action space, favoring memorization over generalizable policies (Chu et al., 2025; Li et al., 2024). While outcome-supervised reinforcement learning (RL) offers a potential solution by exploring reasoning trajectories (Lightman et al., 2023; Paolo et al., 2024), it faces significant hurdles in this domain: the action space in real-world textual graphs is often too vast for efficient exploration, and outcome-based rewards are sparse and noisy. For instance, redundant or erroneous retrieval steps may coincidentally lead to the correct final answer, rendering outcome supervision unreliable for assessing reasoning quality (Liu et al., 2023; Rengarajan et al., 2022).
@@
Extensive experiments have demonstrated the effectiveness of our synthetic stepwise supervision approach. For example, while retrieving only 11.44% of the triples, Graph-S3 achieves an average improvement of 15.6% in accuracy and 17.2% in F1 score across the WebQSP, CWQ, and MetaQA benchmarks.
@@
(2) We design a two-stage training paradigm tailored for graph reasoning: SFT on raw synthetic trajectories to bootstrap basic navigation ability, followed by RL with synthetic stepwise rewards on refined trajectories to provide explicit feedback and strengthen reasoning strategies.
@@ 2.1 Graph Retrieval Methods
Graph retrieval approaches include similarity-based, GNN-based, and LLM-based methods (Peng et al., 2024; Procko and Ochoa, 2024; Zhang et al., 2025a; Zhu et al., 2025; Han et al., 2024; Zhou et al., 2026), but most perform one-shot retrieval and often return redundant or incomplete subgraphs. Recent interactive frameworks (Jiang et al., 2024; Ji et al., 2024; Yang et al., 2024) allow iterative exploration, yet their training predominantly relies on imitation of language patterns or outcome-based supervision, which provides only coarse feedback and limits stable multi-hop reasoning. In contrast, our work employs RL with synthetic stepwise rewards and a scalable data synthesis pipeline to provide supervision for interactive retrieval.
2.2 Stepwise RL for Graph Reasoning
Recent advances such as OpenAI o1 and DeepSeek-R1 (Jaech et al., 2024; Guo et al., 2025) demonstrate the effectiveness of RL in strengthening multi-step reasoning, enabling models to perform longer chains of thought with improved reliability in domains like mathematics and programming (Guo et al., 2024a; El-Kishky et al., 2025; Li et al., 2026; Xu et al., 2025). In contrast, applying RL to textual graphs remains limited, partly due to the lack of fine-grained supervision data (Zhang et al., 2025a; Yao et al., 2025; Liu et al., 2025). RL-based graph agents  (Das et al., 2017; Cui et al., 2025) relied on sparse outcome rewards, making credit assignment across reasoning trajectories difficult. Although recent efforts have introduced reasoning-structured datasets (Pahilajani et al., 2024), high-quality stepwise supervision for graph-based RL remains scarce and difficult to construct at scale. These observations highlight the need for scalable approaches that can provide fine-grained supervision signals and support stable optimization for interactive graph retrieval.
@@ 3 Method
We present our agentic retrieval framework, designed to equip large language models with robust graph reasoning capabilities through stepwise supervision, a two-stage training paradigm, and an interactive retrieval strategy. As illustrated in Figure 2, the framework comprises three main components. First, we construct an automatic data synthesis pipeline that leverages GPT-4o (OpenAI, 2024) to generate diverse exploratory trajectories, which are subsequently refined into high-quality stepwise supervision data. This addresses the scarcity of fine-grained training signals for graph-based RL. Second, we adopt a two-stage training paradigm: SFT on raw synthetic trajectories provides a warm-up initialization for basic graph navigation, while RL with synthetic stepwise rewards on refined trajectories supplies explicit feedback at each decision step, stabilizing optimization and strengthening reasoning strategies. Finally, during inference, the retriever operates under an interactive retrieval mechanism that conducts stepwise, structure-aware exploration of the textual graph, thereby reducing redundancy and mitigating incomplete retrieval.
@@ 3.1.1 Action Space for Graph Exploration
$\text{Explore}(x)=$ $\{(x,r,y)|(x,r,y)\in\mathcal{G}^{p}\}$
$\cup\{(y,r,x)|(y,r,x)\in\mathcal{G}^{p}\}$ (1)
@@
$\begin{split}\operatorname{Choose}(q,\mathcal{G}^{p})=\bigl\{&(x,r,y)\in\mathcal{G}^{p}\big|\\ &F(q,(x,r,y))=1\bigr\}.\end{split}$ (2)
@@ 3.1.2 Graph Reasoning Data Synthesis
Transition. The execution of action $a_{t}$ triggers a state transition $s_{t}\to s_{t+1}$ by updating the action history $h_{t+1}\leftarrow h_{t}\cup\{a_{t}\}$ and modifying the graph components based on the action type. Specifically, if $a_{t}=\textsc{Explore}(x_{t})$, the agent expands the partial graph $\mathcal{G}^{p}_{t}$ by merging it with the newly explored neighbors. If $a_{t}=\textsc{Choose}$, the agent updates the subgraph $\mathcal{G}^{sub}_{t}$ by filtering relevant triples from $\mathcal{G}^{p}_{t}$ via the selection function. The episode terminates when $a_{t}=\textsc{Finish}$ or the time step $t$ reaches the maximum limit $T_{\max}$.
Answer and retention. Upon termination at step $T$, we produce an answer $\hat{y}\;=\;\text{Answer}\!\big(q,\mathcal{G}^{sub}_{T}\big)$ and keep the trajectory $\tau\;=\;\big\{(s_{t},a_{t})\big\}_{t=1}^{T}$ only if $\hat{y}$ matches the set of ground-truth answers. After that, we use the raw action labels for SFT dataset: $\mathcal{D}_{\text{SFT}}\;=\;\big\{(s_{t},a_{t})\big\}_{\tau,\,t}$. In the next subsection, we further refine trajectories to obtain stepwise supervision signals for RL.
3.1.3 Trajectory Refinement for RL
While SFT directly benefits from raw trajectories, RL requires more concise training signals (Yue et al., 2025). Filtering trajectories only by final answer correctness often produces redundant exploration steps, introducing noise and inefficiency during policy optimization. To address this, we introduce a refinement procedure that removes unnecessary detours while preserving all answer-consistent trajectories, thereby yielding shorter and cleaner trajectories for RL.
@@ 3.1.3 Trajectory Refinement for RL
where $\mathcal{F}_{i}$ is the set of all feasible answer-consistent trajectories equivalent to $\tau_{i}$. Let $s_{T}(\tau)$ denotes the final state and $y(\tau)$ denotes the answer.
$\mathcal{F}_{i}=\big\{\tau\mid s_{T}(\tau)=s_{T}(\tau_{i})\land y(\tau)=y(\tau_{i})\big\}.$ (6)
Thus, the refined dataset $\mathcal{T}^{*}$ retains trajectories that are semantically equivalent to the originals but stripped of redundant exploration steps. Each refined trajectory $\tau_{i}^{*}$ is then decomposed into step-level supervision signals by attaching a rule-based stepwise reward $\ell_{t}\in[0,1]$ to each action $a_{t}$, indicating its correctness within the reasoning trajectory. Formally, the RL training dataset is constructed as $\mathcal{D}_{\text{RL}}=\big\{(s_{t},a_{t},\ell_{t})\big\}_{\tau^{*},\,t}$. This ensures that RL receives concise stepwise supervision signals, improving both stability and efficiency of training.
@@ 3.2.1 Stage I: Supervised fine-tuning
$\mathcal{L}_{\mathrm{SFT}}(\theta)=-\mathbb{E}_{\begin{subarray}{c}(s_{t},y_{t})\\ \sim\mathcal{D}_{\text{SFT}}\end{subarray}}\Bigg[\sum_{l=1}^{L_{t}}\log\pi_{\theta}\big(y_{t}^{\,l}\mid\mathcal{I}(s_{t}),\,y_{t}^{<l}\big)\Bigg].$ (7)
@@
3.2.2 Stage II: RL with stepwise rewards
@@
$\ell_{t}=\begin{cases}0,&\text{if }\mathrm{Inv}(a_{t}),\\ c_{1},&\text{if }\mathrm{Fmt}(a_{t})=1\land\mathrm{Corr}(a_{t})=0,\\ c_{2},&\text{if }\mathrm{Part}(a_{t})=1,\\ 1.0,&\text{if }a_{t}=a_{t}^{*}.\end{cases}$ (8)
Here, $\mathrm{Inv}(\cdot)$, $\mathrm{Fmt}(\cdot)$, $\mathrm{Corr}(\cdot)$, and $\mathrm{Part}(\cdot)$ denote the deterministic functions for checking validity, format correctness, action correctness, and partial correctness, respectively. $c_{1}$ and $c_{2}$ are dataset-specific hyperparameters. Consequently, we adopt the reward function shown in Eq. 8 to train our model using the GRPO method.
@@ 3.3 Interactive Retriever
At inference time, Graph-S3 interacts with the textual graph through the defined action space. Unlike single-pass retrieval methods that return large subgraphs, our approach performs stepwise exploration by balancing EXPLORE and CHOOSE actions, terminating with FINISH when sufficient evidence is gathered. This interactive process enables precise control over retrieval depth while minimizing redundancy, producing concise subgraphs for reasoning. The detailed prompts designed for the interactive graph retriever are provided in Appendix B. An example of interactive reasoning process of Graph-$S^{3}$ can be found in Appendix D.
@@ 4 Experiments
Graph-$S^{3}$ + Qwen3-8B 36.24 47.88 17.87 23.29 81.50 90.22 53.73 65.60 12.73 29.49
@@
Graph-$S^{3}$ + Llama3.1-8B 32.31 43.26 17.11 21.17 67.50 76.56 40.17 55.49 10.73 29.55
@@
Graph-$S^{3}$ + Finetuned-8B 44.29 58.45 23.62 30.73 82.77 92.04 63.17 76.18 14.70 36.29